### v1

@@ 4.1 Corpus Collection and Data Preprocessing
Corpus collection was performed by four graduate students from the Department of Sports Science. We first defined inclusion criteria based on common conditions in clinical sports rehabilitation, excluding conditions with scarce literature or lacking authoritative guideline support. Next, we retrieved open-access and institution-subscribed literature on each disease from major databases, such as PubMed and Embase, and supplemented materials from authoritative sports rehabilitation organization websites. Referring to established evidence grading frameworks[46, 47, 48], we categorized the literature types as follows: guidelines and expert consensus, systematic reviews and meta-analyses, RCTs, cohort studies, and other research (case reports, narrative reviews, observational studies, and other corpus). Through targeted search strategies and manual abstract review, we ensured accurate evidence-grade assignment for each document. Ultimately, we compiled a domain corpus covering 21 conditions, mapped to evidence grades A–E.
During data preprocessing, we used Docling to convert PDFs to Markdown and extract hierarchical structure (titles, sections, tables)[49]. We then applied regular expressions to remove header and footer noise (e.g., copyright statements, references) and standardized numeral and unit formatting to reduce cross-document inconsistencies. Finally, we recorded each document’s evidence grade in metadata for subsequent evidence-grade-aware retrieval and BETR calibration.
In the chunking phase, we adopted an LLM-aware hybrid chunking strategy. First, MarkdownHeaderTextSplitter divided documents into sections by heading level[50]. Each section was further split into numbered atomic blocks preserving paragraph structure. Under prompt constraints, the LLM performed semantic grouping of atomic blocks and returned group IDs. The program concatenated each group into final evidence windows. This strategy aimed to enhance the semantic integrity and clinical information density of the windows while preserving paragraph boundaries, aligning with recent LLM-aware chunking methods such as AutoChunker, MetaChunking, and LumberChunker[51, 52, 53].
4.2 BETR Algorithm Workflow
The evidence hierarchy principle in EBM was integrated into the reranking pipeline to optimize ranking order. In existing literature, evidence-hierarchy-based reranking paradigms mostly rely on subjective preset scores, yielding heuristic weighting schemes[41]. Such schemes introduce excessive manually defined rules that reduce interpretability and preclude cross-system reuse. To address this, we proposed BETR. This algorithm introduced evidence hierarchy as an ordered structure into the ranking calibrator, leveraging a data-driven paradigm to improve weight interpretability and cross-system reusability. The complete workflow is presented in Algorithm 1.
@@ 4.2 BETR Algorithm Workflow
2:Calibrator parameters $(\alpha,\delta_{B},\delta_{C},\delta_{D},\delta_{E})$ and online ranking score $r(q,d)$. x
@@
Task Definition and Objective. Given a clinical question $q$ and a candidate evidence window $d$, the reranker outputs an uncalibrated relevance score $s(q,d)$. For the candidate set $\mathcal{C}_{q}$, we aimed to combine the relevance score $s(q,d)$ with evidence hierarchy $\mathrm{Grade}(d)\in\{A\succ B\succ C\succ D\succ E\}$ to yield a unified ranking score $r(q,d)$ satisfying: (1) when semantic relevance differences are large, $s(q,d)$ dominates the ranking; (2) when candidates have comparable $s(q,d)$, higher evidence grades receive higher final scores. This study adopted a unified five-grade evidence hierarchy: A = guidelines and expert consensus, B = systematic reviews and meta-analyses, C = RCTs, D = cohort studies, E = other research (case reports, narrative reviews, observational studies, and other corpus).
Training Labels. For each question $q$, candidate windows included in the reference evidence chain serve as positive examples: $\mathcal{C}_{q}^{+}=\mathcal{C}_{q}\cap\mathcal{W}^{\star}(q)$; the remainder serve as negatives: $\mathcal{C}_{q}^{-}=\mathcal{C}_{q}\setminus\mathcal{W}^{\star}(q)$. This design avoids per-item manual annotation and aligns the ranking objective with evidence window selection.
Pairwise Ranking Objective. BETR adopts a pairwise learning-to-rank approach[54], learning preferences through pairwise comparisons. Specifically, for the candidate pool of question $q$, we construct positive–negative pairs $(d^{+},d^{-})$ with $d^{+}\in\mathcal{C}_{q}\cap\mathcal{W}^{\star}(q)$ and $d^{-}\in\mathcal{C}_{q}\setminus\mathcal{W}^{\star}(q)$, and model preference probability via the Bradley–Terry model[55]:
$P(d^{+}\succ d^{-}\mid q)=\sigma\big(a\Delta s+u_{t^{+}}-u_{t^{-}}\big),$ (3)
@@
Ordered hierarchical parameterization. To explicitly encode a pyramid-shaped evidence hierarchy and prevent grade inversion under noisy labels, we fix the grade ordering as $A\succ B\succ C\succ D\succ E$ and adopt monotonically constrained incremental parameterization for grade effects:
$u_{A}$ $=0,$ (4)

### v2

@@ 3.1 Corpus Collection and Data Preprocessing
We retrieved open-access and institution-subscribed literature from PubMed and Embase, supplemented by authoritative sports rehabilitation organization websites. Referring to established evidence grading frameworks (OCEBM Levels of Evidence Working Group, 2011; National Health and Medical Research Council, 2009; Joanna Briggs Institute, 2014), each document was assigned an evidence grade through manual abstract review: A = guidelines and expert consensus, B = systematic reviews and meta-analyses, C = RCTs, D = cohort studies, E = other research. The final corpus covers 21 common sports rehabilitation conditions. PDFs were converted to Markdown via Docling (Auer et al., 2024), with evidence grades recorded in metadata for downstream retrieval and BETR calibration. For chunking, we adopted an LLM-aware hybrid strategy: documents were first split by heading structure, then the LLM performed semantic grouping of atomic blocks into evidence windows (Jain et al., 2025; Zhao et al., 2024; Duarte et al., 2024); details are provided in Appendix A.
3.2 BETR Algorithm
The evidence hierarchy principle in EBM was integrated into the reranking pipeline to optimize ranking order. Existing evidence-hierarchy-based reranking paradigms mostly rely on subjective preset scores, yielding heuristic weighting schemes (Lu et al., 2025). To address this, we proposed BETR, which introduces evidence hierarchy as an ordered structure into the ranking calibrator via a data-driven paradigm. The complete workflow is presented in Algorithm 1.
@@ 3.2 BETR Algorithm
2:Calibrator parameters $(\alpha,\delta_{B},\delta_{C},\delta_{D},\delta_{E})$ and online ranking score $r(q,d)$.
@@
Task Definition.
Given a clinical question $q$ and a candidate evidence window $d$, the reranker outputs an uncalibrated relevance score $s(q,d)$. We aimed to combine $s(q,d)$ with evidence hierarchy $\mathrm{Grade}(d)\in\{A\succ B\succ C\succ D\succ E\}$ to yield a unified ranking score $r(q,d)$ satisfying: (1) when semantic relevance differences are large, $s(q,d)$ dominates the ranking; (2) when candidates have comparable $s(q,d)$, higher evidence grades receive higher final scores. The five evidence grades follow the hierarchy defined in §3.1.
Training Labels.
For each question $q$, candidate windows included in the reference evidence chain serve as positive examples: $\mathcal{C}_{q}^{+}=\mathcal{C}_{q}\cap\mathcal{W}^{\star}(q)$; the remainder serve as negatives: $\mathcal{C}_{q}^{-}=\mathcal{C}_{q}\setminus\mathcal{W}^{\star}(q)$. This design avoids per-item manual annotation and aligns the ranking objective with evidence window selection. $\mathcal{W}^{\star}(q)$ is constructed via independent relevance grading over the full candidate pool (§3.5) and typically spans multiple evidence grades.
Pairwise Ranking Objective.
BETR adopts a pairwise learning-to-rank approach (Burges et al., 2005), learning preferences through pairwise comparisons. Specifically, for the candidate pool of question $q$, we construct positive–negative pairs $(d^{+},d^{-})$ and model preference probability via the Bradley–Terry model (Bradley and Terry, 1952):
$P(d^{+}\succ d^{-}\mid q)=\sigma\big(a\Delta s+u_{t^{+}}-u_{t^{-}}\big)$ (1)
@@
To explicitly encode a pyramid-shaped evidence hierarchy and prevent grade inversion under noisy labels, we fix the grade ordering as $A\succ B\succ C\succ D\succ E$ and adopt monotonically constrained incremental parameterization:
$u_{A}$ $=0,\quad u_{B}=-\delta_{B},$ (2)
$u_{C}$ $=-(\delta_{B}+\delta_{C}),\quad\ldots,$
$u_{E}$ $=-(\delta_{B}+\delta_{C}+\delta_{D}+\delta_{E})$
where $\delta_{B},\delta_{C},\delta_{D},\delta_{E}\geq 0$, naturally guaranteeing $u_{A}\geq u_{B}\geq u_{C}\geq u_{D}\geq u_{E}$, yielding an evidence-grade pyramid consistent with EBM. We then cast BETR parameter estimation in a Bayesian framework, jointly learning the scale parameter $a$ and grade increments $\boldsymbol{\delta}=(\delta_{B},\delta_{C},\delta_{D},\delta_{E})$ via MAP estimation.
MAP Estimation.
We impose zero-centered Gaussian priors on the parameters, encoding the default assumption of no grade bias: $\alpha\sim\mathcal{N}(0,\sigma_{a}^{2})$, $\delta_{i}\sim\mathcal{N}^{+}(0,\tau^{2})$, where $\mathcal{N}^{+}$ denotes a Gaussian truncated to nonnegative values. This prior has maximum density at $\delta=0$, implying that ranking defaults to being driven by semantic relevance $s(q,d)$; the prior for $\alpha$ is centered at 0, corresponding to $a=\exp(\alpha)\approx 1$, so that semantic scores and grade biases are summed on the same scale. We then fit $(\alpha,\delta_{B},\delta_{C},\delta_{D},\delta_{E})$ via maximum a posteriori estimation (full derivation in Appendix B):
$\max_{\alpha,\;\boldsymbol{\delta}\geq\mathbf{0}}$ $\frac{1}{|\mathcal{Q}_{\mathrm{train}}|}\sum_{q\in\mathcal{Q}_{\mathrm{train}}}\frac{1}{\max(1,|\mathcal{P}_{q}|)}\sum_{(d^{+},d^{-})\in\mathcal{P}_{q}}$ (3)
$\log\sigma\!\big(a\,\Delta s+u_{t^{+}}-u_{t^{-}}\big)$
$-\frac{1}{2\tau^{2}}\big(\delta_{B}^{2}+\delta_{C}^{2}+\delta_{D}^{2}+\delta_{E}^{2}\big)-\frac{1}{2\sigma_{a}^{2}}\alpha^{2}$
where $a=\exp(\alpha)$. The first term is the query-normalized pairwise log-likelihood; the remaining terms are quadratic shrinkage penalties. This framework ensures that the posterior is likelihood-dominated when data are ample and shrinks toward the prior when data are scarce, achieving adaptive regularization. The hyperparameter $\tau$ is selected via grid search on a validation set and fixed across all experiments. At inference, the final ranking score is $r(q,d)=\hat{a}\,s(q,d)+\hat{u}_{\mathrm{Grade}(d)}$ .
3.3 PICO-extended Schema and Knowledge Graph Construction
Schema Definition and Constrained Extraction.
Following Youtu-GraphRAG, the schema is defined as $\mathcal{S}\triangleq\langle\mathcal{S}_{e},\mathcal{S}_{r},\mathcal{S}_{\text{attr}}\rangle$, specifying entity, relation, and attribute types. The LLM extraction agent performs constrained triple extraction: all extracted entity–relation–entity triples and entity–attribute pairs must map to types in $\mathcal{S}$, suppressing out-of-schema noise.
PICO Instantiation.
We instantiate $\mathcal{S}$ as PICO-related types: (1) entity types $\mathcal{S}_{e}^{\text{PICO}}$: Population, Condition, Intervention, Comparator, Outcome, Timepoint, plus domain-specific extensions (e.g., Arm, Device); (2) relation types $\mathcal{S}_{r}^{\text{PICO}}$: directed relations linking studies to PICO entities (e.g., has_population, uses_intervention, reports_outcome); (3) attribute types $\mathcal{S}_{\text{attr}}^{\text{PICO}}$: key dimensions refining PICO elements (e.g., age_bin, followup_weeks, protocol_params). A complete list is provided in Appendix C. The community compression algorithm and knowledge tree indexing follow Youtu-GraphRAG without modification (Appendix C).
3.4 SR-RAG Pipeline
SR-RAG introduces three key improvements to Youtu-GraphRAG: (1) a PICO-guided HyDE channel fused with graph retrieval via RRF; (2) two-stage reranking using ColBERT and a cross-encoder; (3) evidence-grade-aware retrieval followed by BETR-based reranking.
PICO-guided HyDE.
HyDE generates hypothetical documents to bridge the semantic gap between queries and corpus (Gao et al., 2023; Wang et al., 2023). We incorporated PICO soft constraints into the HyDE prompt: available P/I/C/O/T keywords were extracted from the query as anchors (missing fields are permitted). Hypothetical documents must reuse these anchors and are prohibited from fabricating missing fields. This ensures HyDE serves purely as a retrieval intermediary for semantic alignment, without deviating from the original query’s PICO elements.
Two-stage Reranking.
We first used ColBERT (mxbai-edge-colbert-v0) as the coarse ranking model (Takehi et al., 2025). ColBERT uses the MaxSim mechanism for scoring:
where $\mathbf{e}_{i}^{q}$ and $\mathbf{e}_{j}^{d}$ denote token embeddings of the query and window (Khattab and Zaharia, 2020). After coarse ranking, BGE-reranker-v2-m3 serves as a cross-encoder for fine-grained ranking on the top-$K$ candidates (Chen et al., 2024). The two-stage pipeline reduces computational cost while balancing precise term matching and semantic understanding.
Evidence-grade-aware Retrieval.
The corpus was partitioned by evidence grade into Grade A and Grades B–E, forming two candidate pools. Candidate generation and truncation were performed separately on each pool. The candidate sets were then merged, and BETR calibration was applied for final global ranking: $r(q,d)=\hat{a}\,s(q,d)+\hat{u}_{\mathrm{Grade}(d)}$. This design enables the system to prioritize higher-grade evidence when candidate relevance is comparable.
3.5 Benchmark Construction
To facilitate automated evaluation, we created 1,637 QA pairs. Using the annotated corpus, we performed stratified sampling by evidence grade to obtain candidate evidence windows. GPT-4o (OpenAI, 2024) generated a clinical question for each window, strictly grounded in the window’s core conclusion, while simultaneously extracting PICOT elements; inclusion of unsupported information was prohibited (Saad-Falcon et al., 2024). We then applied round-trip consistency filtering (Saad-Falcon et al., 2024): using a system-agnostic hybrid retrieval baseline, questions whose seed window was recalled within top-$K$ entered the main split; the remainder entered a challenge split requiring closer manual inspection.
For reference evidence, we reranked candidates by relevance and had DeepSeek-V3 grade the evidence relationship; strongly supportive and supportive windows formed the gold evidence set. Reference answers were organized by evidence grade, and each answer was decomposed into atomic facts (nuggets) for automated assessment (Min et al., 2023). Four graduate students reviewed all QA pairs from both splits, discarding questions that were clinically irrelevant or lacked practical value, and retaining 1,637 from approximately 2,000 candidates. To validate that LLM-assisted gold labels are reliable, we additionally had two independent annotators re-label gold evidence and PICOT fields from scratch on an 80-query stratified subset (§4.4). The benchmark and knowledge graph are available at https://anonymous.4open.science/r/sr-rag-release-4296/.
3.6 Evaluation
SR-RAG evaluation comprised automated and manual dimensions. For automated evaluation, we used a held-out test set ($n=327$), disjoint from BETR training and validation splits, with five metrics: (1) Evidence recall at 10 (R@10): fraction of gold evidence windows retrieved within the top-10 candidates, a retrieval-level metric independent of LLM judges; (2) Nugget coverage (NC): extent to which the answer covers GT core factual units; (3) Answer faithfulness (Faith.): whether statements are supported by retrieved evidence; (4) Semantic similarity (SS): cosine similarity between answer and GT embeddings; (5) PICOT match accuracy (PM, detailed below).
Mainstream RAG evaluation methods currently lack specific metrics for medical QA, making it difficult to detect PICO mismatches. To address this, we developed PICOT match accuracy (PM). Gold PICOT fields were extracted from reference answers by GPT-4o and spot-checked by annotators; only fields unambiguously stated in the source text were retained. For each system output, PICOT fields are extracted and matched field-by-field against gold fields (synonymous expressions permitted; null gold fields excluded). Let $\mathcal{F}_{q}\subseteq\{P,I,C,O,T\}$ denote the non-null gold fields for query $q$ and $m_{f}\in\{0,1\}$ the match indicator:
$\mathrm{PM}=\frac{1}{N}\sum_{q=1}^{N}\frac{\sum_{f\in\mathcal{F}_{q}}m_{f}}{|\mathcal{F}_{q}|}$ (5)
i.e., the macro-averaged field match rate across queries.
Nugget coverage and PICOT match accuracy were implemented via LLM-as-judge (Zheng et al., 2023). Answer faithfulness and semantic similarity were assessed via the RAGAS framework (Es et al., 2024).
For manual evaluation, five sports rehabilitation experts reviewed 20 randomly sampled questions on a five-point Likert scale across five dimensions: medical factual accuracy, answer faithfulness, answer relevance, safety, and PICOT alignment.
4 Experiments
4.1 Knowledge Graph Construction
Following corpus preprocessing and evidence-grade annotation, we constructed a knowledge graph for sports rehabilitation based on the Youtu-GraphRAG framework. The schema for nodes, relations, and attributes was replaced with PICO-related terms to explicitly encode evidence-based query structure during graph construction (Appendix C). The final knowledge graph consisted of 21 types of sports rehabilitation disease corpora, together with general cross-disease guidelines, comprising a total of 357,844 nodes and 371,226 edges. Of all nodes, 44,033 (12.3%) were core medical entities directly aligned with the PICO framework. Among these, Intervention nodes were most prevalent (13,866; 31.5%), followed by Condition (9,979; 22.7%), Outcome (8,609; 19.5%), and Population (6,838; 15.5%). Arm, Device, and Comparator accounted for 2,134 (4.9%), 1,569 (3.6%), and 1,038 (2.4%), respectively. The graph data will be released upon publication.
Table 1: BETR reranking quality on 327 test queries (K=12K{=}12).
Method AvgGrade HGSR SA-NDCG PrefAcc
Semantic-only 3.18 .362 .781 .523
Heuristic 3.34 .418 .814 .586
BETR 3.51 .473 .847 .641
4.2 BETR Calibration Evaluation
BETR is designed to calibrate ranking so that higher-grade evidence is promoted within the top-$K$ positions, rather than to optimize end-to-end generation quality. We therefore evaluate its reranking behaviour directly. We partitioned the 1,637 benchmark queries into train ($n=983$), validation ($n=327$), and test ($n=327$) splits. BETR parameters were fitted on the train set with hyperparameters selected on the validation set (training details in Appendix B). For each held-out query we retrieved a fixed candidate pool and reranked it under three settings: (1) Semantic-only, ranking by semantic score $s(q,d)$ alone; (2) Heuristic, adding a linearly spaced grade bias $u_{t}^{h}=-c\cdot\mathrm{rank}(t)$ with $c$ tuned on the validation set; (3) BETR, using the learned parameters. We report four ranking-quality metrics at K=12K{=}12 (the operational evidence budget): average evidence grade of gold windows in top-$K$ (AvgGrade), proportion of gold windows that are also high-grade (A+B) among top-$K$ (HGSR, High-Grade Supportive Ratio), support-aware grade NDCG using gold-evidence labels as relevance (SA-NDCG), and gold-vs-non-gold preference accuracy (PrefAcc).
As shown in Table 1, BETR consistently outperforms both baselines across all four metrics, consistent with semantic relevance remaining the primary ranking signal while grade biases provide effective calibration among relevant candidates (the learned scale parameter $a=1.035$ and monotonically decreasing biases are detailed in Appendix B). The gain over Heuristic further demonstrates that data-driven calibration outperforms hand-tuned mappings.
4.3 Baseline Comparison and Ablation Studies
We conducted an end-to-end evaluation of SR-RAG on the 327 held-out test queries, reporting five metrics: evidence recall at 10 (R@10), nugget coverage (NC), answer faithfulness (Faith.), semantic similarity (SS), and PICOT match accuracy (PM). We compared SR-RAG against five baselines and performed ablation experiments on the best-performing model, DeepSeek-V3. All results are shown in Table 2. The benchmark and knowledge graph are available at https://anonymous.4open.science/r/sr-rag-release-4296/.
Baselines.
We compared SR-RAG with five baselines, all using DeepSeek-V3 as the generator and the same corpus: (1) Naive RAG with BM25 + dense retrieval and RRF fusion; (2) Microsoft GraphRAG (Edge et al., 2024) with community-based global and local search; (3) LightRAG (Guo et al., 2025) with dual-level entity–relationship retrieval; (4) Youtu-GraphRAG (Dong et al., 2026), which shares our knowledge graph but uses a generic medical schema without PICO-guided HyDE or BETR; (5) Med-R2 (Lu et al., 2025), an EBM-aligned RAG with a manually defined linear evidence-hierarchy mapping $f_{h}(x){=}9{-}(e_{x}{-}1)$, adapted to our corpus.
Table 2: Automated evaluation results on the SR-RAG benchmark.
[…]
Baselines (DeepSeek-V3)
Naive RAG 0.643 0.718 0.769 0.841 0.582
MS GraphRAG 0.698 0.749 0.784 0.856 0.621
LightRAG 0.721 0.762 0.791 0.862 0.643
[…]
SR-RAG (Ours)
Baichuan-M2 0.812 0.740 0.785 0.806 0.755
[…]
Ablation (DeepSeek-V3)
w/o HyDE 0.764 0.819 0.801 0.879 0.723
w/o ColBERT 0.748 0.798 0.752 0.884 0.740
[…]
w/o BETR 0.810 0.822 0.818 0.881 0.768
Baseline comparison.
SR-RAG (DeepSeek-V3) outperformed all baselines across all five metrics. On the LLM-independent R@10, SR-RAG achieved 0.812 versus 0.643–0.738 for baselines. Youtu-GraphRAG, which shares SR-RAG’s knowledge graph, led baselines on R@10, NC, and SS, confirming the value of graph-structured retrieval. Med-R2, the only EBM-aware baseline, led on Faith and PM, confirming that evidence-hierarchy awareness benefits answer quality; however, SR-RAG still outperformed it on PM by ++16.2% (0.788 vs. 0.678), indicating that learned BETR calibration with PICO-schema-level alignment surpasses manually defined linear mappings.
Model comparison.
DeepSeek-V3 (DeepSeek-AI, 2024) performed best overall, followed by GPT-4o (OpenAI, 2024) (highest faithfulness at 0.842). Although Baichuan-M2 (Baichuan AI, 2025) is medically specialized, its smaller scale limits the complex multi-step reasoning required by SR-RAG.
Ablation analysis.
We ablated four components on DeepSeek-V3 (see Appendix C for schema details). HyDE removal caused the largest R@10 drop (-0.048), confirming its direct retrieval benefit. Removing the PICO schema had the largest impact on NC (-0.056) and PM (-0.087) but slightly increased Faith (+0.015), revealing a coverage–alignment tradeoff that underscores the necessity of domain-specific metrics like PM. ColBERT removal caused the largest Faith drop (-0.067). BETR removal barely affected R@10 (-0.002) but reduced PM by 0.020, consistent with its role as a post-retrieval evidence-hierarchy calibrator rather than a retrieval component.
4.4 Human Evaluation
Expert clinician ratings.
We sampled 20 questions and invited five sports rehabilitation experts for manual evaluation on a 1–5 Likert scale across five dimensions (Table 3; procedures in Appendix E).
Table 3: Expert clinician evaluation results (5 experts, 20 questions, 1–5 Likert scale). Inter-rater SD reflects mean SD across questions per dimension.
Dimension Mean SD Inter-rater SD
Medical factual accuracy 4.71 0.50 0.46
Answer faithfulness 4.84 0.37 0.28
Answer relevance 4.81 0.44 0.33
Safety 4.72 0.57 0.42
PICOT alignment 4.66 0.76 0.63
SR-RAG scored 4.66–4.84 across all dimensions. PICOT alignment showed the largest inter-rater SD (0.63), suggesting sensitivity to question format. To calibrate LLM-based judges, we computed Spearman correlations between automated metrics and corresponding expert dimensions on this 20-question subset: NC vs. expert completeness $\rho=0.68$, Faith. vs. expert accuracy $\rho=0.74$, PM vs. expert PICOT alignment $\rho=0.71$ (all $p<0.01$).
Human-verified subset.
To validate that benchmark quality and system rankings are not artifacts of LLM-assisted gold construction, we stratified-sampled 80 queries from the test set and had two independent annotators re-label gold evidence windows (supportive vs. not) and gold PICOT fields from scratch, with a third annotator adjudicating disagreements ($\kappa_{\text{evidence}}=0.76$, $\alpha_{\text{PICOT}}=0.71$). Table 4 reports R@10 and PM recomputed against these human-verified gold labels.
[…]
Youtu-GraphRAG 0.741 0.718 0.659 0.631
Med-R2 0.724 0.703 0.678 0.649
SR-RAG (DS-V3) 0.812 0.791 0.788 0.762
[…]
5 Conclusion
We present SR-RAG, an EBM-adapted GraphRAG framework for sports rehabilitation. To address the core question of how to adapt the EBM framework to RAG pipelines, we introduce generalizable components organized around two EBM principles: (1) Evidence hierarchy: evidence-grade annotation at the corpus level, BETR for learning grade biases without predefined weights, and dual-track retrieval that recalls guideline evidence separately to prevent dilution; (2) PICO alignment: a PICO-extended graph schema that encodes clinical query structure at the extraction level, and PICO-guided HyDE that bridges the query–evidence semantic gap with soft PICOT constraints. We also construct the first knowledge graph (357,844 nodes, 371,226 edges) and benchmark (1,637 QA pairs) for sports rehabilitation. Automated evaluation, expert clinician review, and a human-verified gold subset all confirm SR-RAG’s reliability.
Limitations
First, expert evaluation is limited in scale (20 questions, 5 experts). Second, BETR operates at the study-type level without assessing within-study evidence quality (e.g., risk of bias); integrating fine-grained quality dimensions such as GRADE (Wang et al., 2025) into reranking is a promising direction.
Appendix A Chunking Strategy
In the chunking phase, we adopted an LLM-aware hybrid chunking strategy. First, MarkdownHeaderTextSplitter divided documents into sections by heading level (Chase, 2022). Each section was further split into numbered atomic blocks preserving paragraph structure. Under prompt constraints, the LLM performed semantic grouping of atomic blocks and returned group IDs. The program concatenated each group into final evidence windows. This strategy aimed to enhance the semantic integrity and clinical information density of the windows while preserving paragraph boundaries, aligning with recent LLM-aware chunking methods such as AutoChunker, MetaChunking, and LumberChunker (Jain et al., 2025; Zhao et al., 2024; Duarte et al., 2024).
Appendix B BETR Full Derivation
B.1 Training Labels
For each question $q$, candidate windows included in the reference evidence chain serve as positive examples: $\mathcal{C}_{q}^{+}=\mathcal{C}_{q}\cap\mathcal{W}^{\star}(q)$; the remainder serve as negatives: $\mathcal{C}_{q}^{-}=\mathcal{C}_{q}\setminus\mathcal{W}^{\star}(q)$. This design avoids per-item manual annotation and aligns the ranking objective with evidence window selection.
B.2 Pairwise Ranking Objective
BETR adopts a pairwise learning-to-rank approach (Burges et al., 2005), learning preferences through pairwise comparisons. Specifically, for the candidate pool of question $q$, we construct positive–negative pairs $(d^{+},d^{-})$ with $d^{+}\in\mathcal{C}_{q}\cap\mathcal{W}^{\star}(q)$ and $d^{-}\in\mathcal{C}_{q}\setminus\mathcal{W}^{\star}(q)$, and model preference probability via the Bradley–Terry model (Bradley and Terry, 1952):
$P(d^{+}\succ d^{-}\mid q)=\sigma\big(a\Delta s+u_{t^{+}}-u_{t^{-}}\big),$ (B.1)
B.3 Ordered Hierarchical Parameterization
To explicitly encode a pyramid-shaped evidence hierarchy and prevent grade inversion under noisy labels, we fix the grade ordering as $A\succ B\succ C\succ D\succ E$ and adopt monotonically constrained incremental parameterization:
$u_{A}$ $=0,$ (B.2)